a car starts from rest and accelerates uniform rate of 5 m/s^2 for sometime and retards at the same rate and comes to rest. In between acceleration and retardation it also moves with uniform velocity.Total time for journey is 25 sec and average speed for journey is 20 m/s . How long does the car moves with constant speed

Grade:11

1 Answers

Vikas TU
7 years ago
During acceleration,
x1 = 0 + 0.5*5*t1^2..............(1)
uniform motion,
x2 = v*t2.............(2)
retardation,
x3 = v*t3 – 0.5*5*t3^2..........(3)
t1 + t2 + t3 = 25 …................(4)
avg. speed,
20 = (x1 + x2 + x3)/25.....................(5)
Just before uniform motion,
v = 0 + 5*t1.......................(6)
solve the eqns. and get the values.